OpenStudy (anonymous):

The lengths of the triangle are 6,8 and 12. if the length of the shortest side of a triangle is 15, what is the perimeter of the larger triangle?

OpenStudy (anonymous):

If the shortest side of a similar triangle is 15 take 15 and divide it by 6 which equals 2.5 then find larger values that when divided by 8&10 will equal 2.5 15/6=2.5 20/8=2.5 25/10=2.5 So the values of the triangle are 15,20,25. Added together they equal 60 and the smaller triangle equals 24 60/24=2.5 so the similarity ratio for the triangles would be 2.5:1 which is the largest triangle: to the smallest or 1:2.5 which is the smallest triangle: to the larger one. Values of the triangle: 15,20,25 Ratio: 2.5:1 or 1:2.5 (Replace 10 with 12) hope tis helped

OpenStudy (anonymous):

Thank Ya Sugga! :)

OpenStudy (anonymous):

not a problem at all(:

OpenStudy (anonymous):

But one question... what was the purpose of replacing 12 with 10? :)

OpenStudy (anonymous):

because you have 6,8, and 12 not 6,8, and 10 am I right?

OpenStudy (anonymous):

either way I guess it wouldn't matter

OpenStudy (anonymous):

Correct. So in order for me to find the perimeter of the larger triangle, i would have to get 2.5 when i divide my measures given (6,8 and 12) right?

OpenStudy (anonymous):

15/6=2.5 20/8=2.5 25/10=2.5. yes you are correct

OpenStudy (anonymous):

Okay, Thanks again.

OpenStudy (anonymous):

not a problem(: